Steve hit four more home runs than twice the number of home runs Larry hit. Together they hit 10 home runs. What is the equation for finding how many home runs Steve hit?
Answer by checkley75(3666)

(2x+4)+x=10
2x+4+x=10
3x=10-4
3x=6
x=6/3
x=2 number of home runs Larry hit.
2*2+4=4+4=8 numbe of home runs Steve hit.
Proof
8+2=10
10=10